OpenStudy (anonymous):

PLEASE HELP!! Simplify using i. √-15 √-4

myininaya (myininaya):

\[\sqrt{-1}=i\] you have this: \[\sqrt{-1} \sqrt{-1} \sqrt{15}\sqrt{4}\]

myininaya (myininaya):

replace the sqrt(-1)s with i

myininaya (myininaya):

and you can do sqrt(4) I'm sure.

OpenStudy (anonymous):

so would it be i sqrt 15 and i sqrt 4?

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

well i*i is not i

myininaya (myininaya):

do you know what i^2 equals?

OpenStudy (jdoe0001):

\(\bf \sqrt{-15}\cdot \sqrt{-4}\implies \sqrt{-1\cdot 15}\cdot \sqrt{-1\cdot 4}\implies \sqrt{-1}\cdot \sqrt{15}\cdot \sqrt{-1}\cdot \sqrt{4} \\ \quad \\ \sqrt{-1}\cdot\sqrt{-1}\cdot \sqrt{15}\cdot \sqrt{4}\implies i\cdot i\cdot \sqrt{15\cdot 4}\)

OpenStudy (anonymous):

sqrt -2 right?

myininaya (myininaya):

nope

OpenStudy (jdoe0001):

\(\bf i^2\implies \sqrt{-1}\cdot\sqrt{-1}\implies \sqrt{(-1)^2}\implies ?\)

myininaya (myininaya):

this should be something your teacher told you I'm sure think back to the lecture your teacher most likely said what i^2 is equal to

OpenStudy (anonymous):

ohh -1

myininaya (myininaya):

ok and what is sqrt(4)

myininaya (myininaya):

\[i*i \sqrt{4} \sqrt{15}=-1 \sqrt{4} \sqrt{15} \]

myininaya (myininaya):

what integer numbers times itself will give you 4?

OpenStudy (anonymous):

2

myininaya (myininaya):

so sqrt(4)=2 so you have \[-1(2) \sqrt{15}=-2 \sqrt{15}\]

myininaya (myininaya):

i wrote -2 (in front of sqrt(15) instead of -1(2) since -1 times 2 is -2

OpenStudy (anonymous):

Ohh okay, I think I get it now, Thanks @myininaya

myininaya (myininaya):

np do you want to try one on your own and I check it?

myininaya (myininaya):

\[\sqrt{-9} \sqrt{-16}\]